solve each equation
x/3=4/x+1

Found 2 solutions by mananth, unlockmath:
Answer by mananth(16946)

x/3=4/x+1
x/3 = 4/ (x+1)
x(x+1)=12
x^2+x-12=0
x^2+4x-3x-12=0
x(x+4)-3(x+4)=0
(x-3)(x+4)=0
x=3 OR -4

Answer by unlockmath(1688)

Hello,
With this x/3=4/x+1 let's multiply everything by 3x to get:(Note: I'm assuming the problem is written as x/3=(x/4)+1
x^2=12+3x Subtract 12 and 3x from both sides to get:
x^2-3x-12=0 Since this doesn't factor cleanly, we'll need to us the quadratic formula.
Here's the formula x = (-b +- sqrt( b^2-4*a*c ))/(2*a)
I'll let you complete the problem.
RJ
www.math-unlock.com